# Why Do Open-Source LLMs Struggle with Data Analysis? A Systematic Empirical Study

### v2

@@ Abstract
Large Language Models (LLMs) hold promise in automating data analysis tasks, yet open-source models face significant limitations in these kinds of reasoning-intensive scenarios. In this work, we investigate strategies to enhance the data analysis capabilities of open-source LLMs. By curating a seed dataset of diverse, realistic scenarios, we evaluate models across three dimensions: data understanding, code generation, and strategic planning. Our analysis reveals three key findings: (1) Strategic planning quality serves as the primary determinant of model performance; (2) Interaction design and task complexity significantly influence reasoning capabilities; (3) Data quality demonstrates a greater impact than diversity in achieving optimal performance. We leverage these insights to develop a data synthesis methodology, demonstrating significant improvements in open-source LLMs’ analytical reasoning capabilities.
Why Do Open-Source LLMs Struggle with Data Analysis?
A Systematic Empirical Study
Yuqi Zhu♠♡, Yi Zhong, Jintian Zhang♠♡, Ziheng Zhang, Shuofei Qiao♠♡, Yujie Luo♠♡, Lun Du♣♡, Da Zheng♣♡, Huajun Chen♠♡, Ningyu Zhang♠♡thanks:     Corresponding Author. Zhejiang University  Ant Group  Independent Researcher Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph {zhuyuqi,zhangningyu}@zju.edu.cn
1 Introdution
@@
Data analysis is a complex and multifaceted task that lies at the core of many real-world applications Donoho (2017); Inala et al. (2024a). It involves a series of interdependent steps, including understanding user queries, exploring and interpreting structured data, formulating hypotheses, and generating insights. As shown in Figure 1, unlike traditional tasks, data analysis requires tight integration of natural language understanding, logical reasoning, code drafting, and result interpretation—all within an interactive and goal-driven setting.
Large Language Models (LLMs) have emerged as promising tools to automate data analysis tasks, especially given their strong capabilities in reasoning, planning, and code generation. To enhance their performance on these complex tasks, researchers have explored strategies such as prompting techniques, fine-tuning, and multi-agent frameworks Guo et al. (2024); Trirat et al. (2024); Hu et al. (2024); Hong et al. (2024). Additionally, domain-specific datasets like DSBench Jing et al. (2024) and BLADE Gu et al. (2024) have been developed to adapt models to these unique challenges.
Despite the promise, current state-of-the-art performance is often driven by large-scale, highly advanced models (e.g., GPT-4 Hurst et al. (2024), DeepSeek-R1 DeepSeek-AI et al. (2025)), which can limit accessibility and pose challenges for advancing research. Open-source small models, while promising, continue to lag in capability, particularly for reasoning-intensive applications. This raises a key research question: How can we effectively enhance open-source LLMs for complex, reasoning-intensive data analysis tasks? Inspired by prior work on fine-tuning with high-quality synthetic data Muennighoff et al. (2025); Ye et al. (2025); Ahmad et al. (2025), we explore whether similar strategies can be leveraged to improve open-source LLMs in the data analysis domain.
@@
To answer the above question, we conduct a systematic exploration of the essential capabilities required for data analysis and propose methods to develop these capabilities in open-source LLMs. Specifically, we use Qwen models as the foundational baseline and evaluate their performance across three critical dimensions: Data Comprehension, Code generation, and Strategic Planning. Leveraging a curated seed dataset that encompasses diverse data analysis scenarios, we conduct targeted experiments and ablation studies to analyze the factors influencing model generalization and performance on complex analytical tasks.
Our analysis reveals three key findings: (1) The model’s planning ability emerges as a more critical determinant of success than its capabilities in data understanding and code generation. This underscores the importance of strategic foresight and structured reasoning in navigating complex analytical scenarios. (2) Appropriate interaction turns, combined with the complexity of the data and suitable reasoning descriptions, can enhance the model’s reasoning capacity. However, performance gains vary across different data analysis tasks, suggesting that task-specific characteristics play a crucial role in shaping the model’s effectiveness. (3) High-quality training data proves to be more critical than data diversity for achieving optimal performance in data analysis tasks. This emphasizes the necessity of curating datasets with precise and comprehensive annotations to ensure reliable and robust model outcomes. To further substantiate these insights, we synthesize new training data informed by the identified strategies and subsequently retrain the model, thereby validating the effectiveness of our approach.
@@ 2 Background of Data Analysis Agents
Data analysis tasks aim to derive actionable insights from data through systematic exploration and analysis Inala et al. (2024b); Sun et al. (2024). In a typical workflow, analysts begin with specific questions about a dataset, then proceed through multiple analytical steps. These steps include data preprocessing and cleaning, hypothesis exploration, data transformations, and report generation. The process is inherently interactive - analysts work with structured tabular data, develop analysis code, interpret intermediate results, and iteratively refine their analysis before generating final reports with their findings.
To formally characterize the data analysis process in the context of LLM agents, we define a parameterized analysis function $f_{\theta}$ that maps the input components to the analytical outputs:
@@
3 Dataset and Experiments
To investigate the capabilities of different models in the domain of data analysis, we first collect a dataset tailored to this field and introduce the experimental setup in this section.
3.1 Data Collection and Curation
Initial Collection.
To construct our training dataset, we gather samples from DAEval Hu et al. (2024), which provides real-world CSV files from GitHub repositories, DSBench Jing et al. (2024) containing analysis tasks from ModelOff competitions, and TableBench Wu et al. (2025) featuring complex tabular tasks across multiple domains. We further include structured data from WTQ Pasupat and Liang (2015) and FetaQA Nan et al. (2022). To investigate the impact of different chain-of-thought reasoning strategies on data analysis tasks, we employ DeepSeek-R1 to generate synthetic datasets tailored for this study. All training data are collected to ensure no overlap with our evaluation benchmarks, enabling a reliable out-of-distribution (OOD) assessment. Utilizing correctness-based filtering, we curated 6,443 distinct samples, encompassing a diverse array of analytical challenges (refer to Appendix B for detailed statistics).
Filtering Criteria.
To ensure data quality, we apply a two-stage filtering process. First, we automatically remove invalid samples based on several criteria: low-quality code implementations, such as those failing to utilize provided files or producing code with no meaningful return values; samples containing compilation errors; and entries that do not conform to format requirements. Following the automated filtering, we perform manual verification through sampling to further refine the dataset. This process results in a final collection of 5,613 high-quality samples for subsequent experiments.
3.2 Experiment Setup.
@@ 3.1 Data Collection and Curation
We evaluate a series of open-source model variants, including Qwen2.5-7B-Instruct, Qwen2.5-7B-Coder, Qwen2.5-7B-Instruct-1M, and R1-Distill-Qwen-7B, alongside strong baselines like GPT-4o, DeepSeek-v3111We use gpt-4o-2024-08-06 and deepseek-v3-0324., and DeepSeek-R1. Following prior work (Hu et al., 2024; Zeng et al., 2025), we adopt ReAct (Yao et al., 2023) to iteratively alternate between planning, code generation, and execution for multi-turn data analysis. During the subsequent experimental analysis, we adopt two approaches: prompt-based analysis and fine-tuning using Low-Rank Adaptation (LoRA) Hu et al. (2022). Model training is based on Qwen2.5-Instruct, with hyperparameters, experimental configurations, and their rationale detailed in Appendix A.
Evaluation.
@@
To assess models’ performance on data analysis tasks, we employ two comprehensive benchmarks: DiscoveryBench Majumder et al. (2024) and QRData (Liu et al., 2024). DiscoveryBench covers 264 tasks across six domains (e.g., sociology, engineering), and we focus on its real-world test set comprising 239 tasks. QRData, specifically designed for statistical and causal analysis, consists of 411 questions paired with data sheets sourced from textbooks, materials, online learning and academic papers. Following prior work Zeng et al. (2025), we adopt accuracy as our evaluation metric. Since both model outputs and ground truth are expressed in natural language, we employ GPT-4o-mini222gpt-4o-mini-2024-07-18 for agreement evaluation.
@@ 4 Core Capabilities for Data Analysis
Data analysis tasks present a unique challenge, requiring the integration of multiple capabilities. The process encompasses three essential skills that determine model performance: data comprehension, which addresses the ability to understand and utilize structured data effectively; code generation , which focuses on writing correct and efficient analytical solutions through programming; and strategic planning , which involves understanding problems and breaking them down into manageable steps. These three aspects - Data comprehension (§4.1), Code generation (§4.2) and Strategic planning (§4.3) - form the foundation of our investigation into data requirements for effective analytical capabilities.
@@
Table 1: Performance comparison across models in multi-turn settings.
4.1 Data Comprehension
Tabular Information.
To evaluate the model’s ability to understand tabular data, we provide both the table content and the associated reasoning path as input. This setup allows us to examine whether explicitly including table information enhances the model’s reasoning performance. As shown in Table 2, adding table information helps the models in simpler tasks like QRData, showing slight improvements in performance. However, the gains are limited, indicating that the models already handle much of the reasoning without explicit table inputs. For the more complex DiscoveryBench, while the 7B model benefits from the inclusion of table information, the 14B model exhibits a slight drop in accuracy. This decline can be attributed to the increased input length introduced by the table input, which leads the 14B model to generate longer output sequences. The expanded output length appears to negatively impact the model’s reasoning performance in this specific case.
Model QRData DiscoveryBench
w/o Info w Info w/o Info w Info
Qwen2.5-7B 6.57 7.54 0.42 1.26
Qwen2.5-14B 15.09 15.82 0.42 0.00
Data Complexity.
To further evaluate the robustness of the model under increasing data complexity, we introduce additional tables as redundant information, including distractors unrelated to the task. This setup requires the model to reason over multiple tables, some of which are irrelevant, simulating scenarios with heightened complexity. Models that effectively filter out irrelevant data while focusing on task-relevant information are considered to exhibit stronger data understanding capabilities. For a fair evaluation, we remove task-specific background descriptions from the input, forcing the model to rely solely on the tabular data for reasoning. As shown in Table 3, the inclusion of redundant data increases input complexity but does not lead to a significant decline in overall performance. The 7B model exhibits a modest decline in accuracy on both benchmarks, suggesting it is more sensitive to increased input noise. In contrast, the 14B model demonstrates stronger robustness, maintaining stable performance across both datasets.
Model QRData DiscoveryBench
w/o Extra w/ Extra w/o Extra w/ Extra
Qwen2.5-7B 37.96 34.55 5.44 4.18
Qwen2.5-14B 52.55 52.07 10.88 12.13
[Uncaptioned image] Discussion and Implications.
Data information exerts limited influence on models, particularly larger ones (e.g., 14B), which maintain stable performance even as input complexity increases. This suggests that the ability to process structured data is not the primary limitation for larger models; instead, their performance may be constrained by other underlying bottlenecks.
@@ 4.2 Code Capability
To explore the relationship between code generation capabilities and data analysis performance, we conduct a series of experiments using models that vary in scale and specialization.
Code Performance.
Our evaluation here examines whether code-specialized pretraining improves task performance and explores the trade-offs between code generation accuracy and model efficiency in analytical tasks. Table 1 provides an overview of the overall model perfomance, while Table 4 summarizes the average error rates observed across two datasets.
Our analysis reveals several key findings: (1) Limited impact of code specialization: Qwen2.5-7B-Coder does not demonstrate a clear advantage over general-purpose models. This suggests that code specialization alone may not directly translate to better performance in analytical tasks, possibly due to limitations in instruction-following or reasoning generalization. (2) Hallucination in distilled models: R1-Distill-Qwen-7B, despite being distilled from a long chain of thought model, performs poorly, often hallucinating file interpretations rather than generating executable code. (3) Efficiency in task execution: When comparing Qwen2.5-7B-Instruct-1M and Qwen2.5-7B-Instruct under matched output length constraints, both models exhibit comparable coding capabilities; however, the latter demonstrated superior planning efficiency by completing tasks in fewer interaction rounds.
To further substantiate these findings, we manually sample 354 erroneous responses generated by the and categorized the errors using GPT-4o-mini. The categorization is based on the gap between the incorrect responses and the corresponding correct trajectories. As shown in Figure 2, only a small fraction of the errors are attributed to deficiencies in code generation capabilities. Instead, the majority of errors are linked to shortcomings in the models’ planning and reasoning abilities.
@@
[Uncaptioned image] Discussion and Implications.
The ability to generate code alone is insufficient to guarantee strong performance in data analysis tasks. Instead, a model’s capacity to effectively plan and seamlessly integrate code generation with reasoning emerges as a more decisive factor.
@@ 4.3 Strategic Planning
Data analysis is inherently a planning-intensive process that demands careful coordination of data access, transformation, and reasoning steps. Building on prior experiments, we further investigate how the strategic reasoning capabilities of LLMs influence data analysis. Specifically, we evaluate their performance along four key dimensions: Interaction Turns, Reasoning Length, Task Complexity and Problem Diversity.
4.3.1 Interaction Turns
@@
Trend across Different Model Sizes.
Figure 3 reports the performance of Qwen2.5-7B and Qwen2.5-14B under different dialogue strategies. To ensure a fair comparison, both models were fine-tuned using the same dataset size across all strategies, which is 1020 here. The results reveal consistent trends across the two models: medium-length interactions generally achieve relatively better performance across both datasets, suggesting their effectiveness for tasks requiring moderate reasoning depth. In contrast, short and long interaction strategies yield slightly lower yet relatively stable results. However, the mixed strategy consistently demonstrates inferior performance, potentially due to the variation in turn lengths, which may hinder the model’s ability to capture consistent interaction patterns during training. Then given the alignment in performance trends between the two models, we selected the 7B model for subsequent experiments. This decision balances computational efficiency with analytical rigor, enabling deeper exploration of dialogue strategies.
@@
Table 5: Effectiveness analysis on different turn lengths.
Further Investigation.
To better understand the impact of turn length, we examine the performance of various interaction strategies using the collected dataset. The results are presented in Table 5, from which we derive the following observations:
(i) Turn length preferences differ across tasks.
@@
(iii) Mixed strategies demand thoughtful optimization.
Mixed strategies, such as Medium + Short and Medium + Long, exhibit balanced but slightly lower performance. These observations suggest that different interaction styles may offer complementary strengths, but leveraging them effectively likely requires a carefully designed curriculum or selection mechanism rather than uniform mixing.
@@
4.3.2 Reasoning Length
To investigate whether longer reasoning chains from stronger models improve planning and task success, we augment training samples with intermediate <think> segments generated by DeepSeek-R1. These segments aim to capture explicit intermediate reasoning steps that may scaffold better decision-making during multistep analysis.
We evaluate three settings: (1) Original reasoning, which uses the original training data without modification. These samples include reasoning content, but the reasoning is typically shorter and less explicit compared to the augmented settings; (2) Full reasoning, which replaces the original reasoning with full <think> traces for intermediate turns (excluding the first and final turns); and (3) Summarized reasoning, which substitutes the original reasoning with concise summaries (generated by an LLM) of the full traces. The reasoning is inserted only in the middle turns to avoid hallucinations during initial grounding and to maintain brevity in the final answer generation. Experiments are conducted under varying per-turn token budgets (1024, 2048, and 4096) to simulate different interaction constraints. Results are reported in Figure 4 from which we get the following findings:
(i) Longer are not always better.
While longer reasoning chains might seem beneficial, the Full setting consistently underperforms the Original across most configurations, particularly in DiscoveryBench. The sharp performance decline at the 4096-token level suggests that overly verbose and unfiltered reasoning may overwhelm the model. In contrast, the Summarized reasoning setting, which provides more detailed reasoning than the Original while avoiding the verbosity of the Full, consistently matches or surpasses baseline accuracy. By striking a balance between informativeness and brevity, summarized reasoning preserves essential steps and enhances clarity, offering a more effective approach to robust decision-making. These findings indicate that longer reasoning chains are not inherently advantageous, and appropriately detailed yet concise reasoning may lead to better outcomes.
@@ 4.3.2 Reasoning Length
Increasing the per-turn token budget can improve performance by enabling better integration of reasoning and task-relevant content, as seen in datasets like QRData. However, these gains are often limited, and in some cases, such as with DiscoveryBench, a larger token budget may even reduce performance by amplifying noise or irrelevant information. This highlights that more tokens are not always better; the effectiveness of token budgets depends on the specific task and dataset. Concise and high-quality reasoning is often more effective than simply increasing the token budget.
Figure 4: Performance comparison of QRData and DiscoveryBench across different settings and token limits.
@@
Table 6: Analysis of Complexity characteristics using optimal turn settings (accuracy). Each difficulty level contains 733 samples.
4.3.3 Task Complexity
To evaluate the impact of data difficulty on model reasoning capabilities, we classify each example based on the performance of models with varying capacities. Specifically, a task is labeled easy if it can be correctly solved by Qwen2.5-7B, medium if only Qwen2.5-14B can solve it, and hard if it requires DeepSeek-R1 to provide the correct answer.
We train the model using data of varying difficulty levels, with the results summarized in Table 6. Analysis of the results reveals a clear trend: the model’s overall performance improves as the difficulty of the training data increases.
To further investigate the impact of task complexity on the model’s behavior, we analyze the models’ dialogue turns and average response length, as illustrated in Figure 5. The findings show that: As the difficulty of training data increases, the model’s responses shift from relying on interactions with more turns to obtain feedback to providing more comprehensive answers in fewer turns. This also implies an increase in the workload per step for the model and improvement in model capabilities.
From the results in Table 6, the comprehensive effects of medium and hard data are better, especially on QRData, which is relatively simple. Notably, when medium-level and hard-level data are combined during training, the model achieves the lowest average number of interaction turns and the shortest average response length, indicating that this mixed-difficulty strategy cultivates diverse problem-solving approaches and maximizes answer efficiency. However, on the more complex DiscoveryBench dataset, models trained on simple data performed better, likely due to the dataset’s need for more dialogue turns—a behavior aligned with models trained on simpler tasks.
Figure 5: Analysis of Difficulty. (a) Average number of response rounds of the model. (b) Average output token length of the model.
4.3.4 Problem Diversity

### v3

@@ Abstract
Large Language Models (LLMs) hold promise in automating data analysis tasks, yet open-source models face significant limitations in these kinds of reasoning-intensive scenarios. In this work, we investigate strategies to enhance the data analysis capabilities of open-source LLMs. By curating a seed dataset of diverse, realistic scenarios, we evaluate model behavior across three core dimensions: data understanding, code generation, and strategic planning. Our analysis reveals three key findings: (1) Strategic planning quality serves as the primary determinant of model performance; (2) Interaction design and task complexity significantly influence reasoning capabilities; (3) Data quality demonstrates a greater impact than diversity in achieving optimal performance. We leverage these insights to develop a data synthesis methodology, demonstrating significant improvements in open-source LLMs’ analytical reasoning capabilities.
1 Introduction
@@
Data analysis is a complex, interactive process central to scientific discovery, business intelligence, and decision-making (Donoho 2017; Inala et al. 2024a). It requires models to understand natural language queries, interpret structured data, formulate hypotheses, generate executable code, and iteratively refine reasoning—often across multiple turns of interaction (Figure 1). As such, it demands tight integration of language understanding, logical reasoning, programming skills, and long-horizon planning, which together pose unique challenges beyond standard NLP tasks.
Large Language Models (LLMs) have shown promise in automating data analysis, with systems like DS-Agent (Guo et al. 2024), AutoML-Agent (Trirat, Jeong, and Hwang 2024), and Data Interpreter (Hong et al. 2024) demonstrating increasingly sophisticated behaviors. Domain-specific benchmarks such as DSBench (Jing et al. 2024), BLADE (Gu et al. 2024), QRData (Liu et al. 2024), and DiscoveryBench (Majumder et al. 2024) further enable targeted evaluation. Yet, performance on these tasks remains dominated by large-scale, advanced models such as GPT-4 (Hurst et al. 2024) and DeepSeek-R1 (DeepSeek-AI et al. 2025), while open-source alternatives, especially smaller models, still struggle in real-world analytical settings.
This raises a key research question: How can we effectively enhance open-source LLMs for complex, reasoning-intensive data analysis tasks? Prior work has shown that fine-tuning on high-quality synthetic data can improve reasoning capabilities in domains like mathematics (Muennighoff et al. 2025; Ye et al. 2025) and code generation (Ahmad et al. 2025). However, for data analysis tasks that involve multi-step interactions, dynamic environments, and mixed goals, it remains unclear which properties of the training data, such as task difficulty, scenario diversity, or interaction structure, actually lead to better generalization.
@@
In this paper, we take a capability-aware approach to understanding how to build more effective data analysis agents by analyzing the model capabilities involved in the analytical process. We decompose the task into three core dimensions: Data Comprehension, Code Generation, and Strategic Planning. Leveraging a curated seed dataset that encompasses diverse data analysis scenarios, we conduct targeted experiments and ablation studies to analyze the factors influencing model generalization and performance on complex analytical tasks. Here, our analysis reveals three key findings:
  • The model’s planning ability emerges as a more critical determinant of success than its capabilities in data understanding and code generation. This underscores the importance of strategic foresight and structured reasoning in navigating complex analytical scenarios.
  • Appropriate interaction turns, combined with the complexity of the data and suitable reasoning descriptions, can enhance the model’s reasoning capacity. However, performance gains vary across different data analysis tasks, suggesting that task-specific characteristics play a crucial role in shaping the model’s effectiveness.
  • High-quality training data proves to be more critical than data diversity for achieving optimal performance in data analysis tasks. This emphasizes the necessity of curating datasets with precise and comprehensive annotations to ensure reliable and robust model outcomes.
To substantiate these insights, we propose a strategy-guided data synthesis framework that leverages empirical findings to inform key design choices, such as selecting informative interaction patterns and enriching data with concise reasoning traces, to enable more effective model learning. By fine-tuning on the resulting dataset, we demonstrate measurable improvements in performance, achieving results competitive with leading closed-source models.
@@ 2 Background of Data Analysis Agents
Data analysis tasks aim to derive actionable insights from data through systematic exploration and analysis (Inala et al. 2024b; Sun et al. 2024). In a typical workflow, analysts begin with specific questions about a dataset, then proceed through multiple analytical steps. These steps include data preprocessing and cleaning, hypothesis exploration, data transformations, and report generation. The process is inherently interactive - analysts work with structured tabular data, develop analysis code, interpret intermediate results, and iteratively refine their analysis before generating final reports with their findings.
To formally characterize the data analysis process for LLM agents, we define a parameterized analysis function $f_{\theta}$ that maps the input components to the analytical outputs:
@@
3 Experimental Settings
Evaluation Protocol
We adopt a capability-aware evaluation protocol aligned with our decomposition into Data Comprehension, Code Generation, and Strategic Planning. For the former two, we use prompt-based evaluation; for Strategic Planning, we employ LoRA-based fine-tuning (Hu et al. 2022) to target long-horizon reasoning.
Model Selection
@@ 3 Experimental Settings
We evaluate a series of open-source model variants, including Qwen2.5-7B-Instruct, Qwen2.5-7B-Coder, Qwen2.5-7B-Instruct-1M, and R1-Distill-Qwen-7B, alongside strong API-based baselines like GPT-4o, DeepSeek-v3111gpt-4o-2024-08-06, deepseek-v3-0324, and DeepSeek-R1. All models follow the ReAct framework (Yao et al. 2023) for multi-turn interaction, iteratively alternating between planning, code generation, and execution.
Data Collection and Curation
To construct the training dataset, we collect samples from DAEval (Hu et al. 2024) (real-world CSVs from GitHub), DSBench (Jing et al. 2024) (ModelOff competition tasks), TableBench (Wu et al. 2025) (multi-domain tabular reasoning), WTQ (Pasupat and Liang 2015), and FetaQA (Nan et al. 2022). To enrich reasoning diversity, we generate additional synthetic samples all using DeepSeek-R1.
All training data are collected to ensure no overlap with our evaluation benchmarks, enabling a reliable out-of-distribution (OOD) assessment. Utilizing correctness-based filtering, we retain 6,443 distinct samples, encompassing a diverse array of analytical challenges. To ensure data quality, we apply a two-stage filtering process. First, we remove invalid samples based on several criteria: low-quality code implementations, such as those failing to utilize provided files or producing code with no meaningful return values; samples containing compilation errors; and entries that do not conform to format requirements. Following the automated filtering, we perform manual verification through sampling to further refine the dataset. The final corpus contains 5,613 high-quality instances used for fine-tuning.
Evaluation
@@
We evaluate on two comprehensive benchmarks: DiscoveryBench (Majumder et al. 2024), which includes 239 real-world tasks from its 264-task suite across six domains (e.g., sociology, engineering), and QRData (Liu et al. 2024), a benchmark designed for statistical and causal analysis comprising 411 questions paired with data sheets from textbooks, academic papers, and online resources.
Following Zeng et al. (2025), we use accuracy as the evaluation metric. Given that both predictions and references are in natural language, we employ GPT-4o-mini222gpt-4o-mini-2024-07-18 for agreement scoring. Further details on dataset statistics and training configurations are provided in Appendix A.
@@ 4 Core Capabilities for Data Analysis
Data analysis tasks present a unique challenge, requiring the integration of multiple capabilities. We identify three core skills that determine model performance: Data Comprehension, the ability to understand and effectively utilize structured data; Code Generation, the skill of producing correct and efficient analytical code; and Strategic Planning, the capacity to decompose complex problems into manageable steps. All ablation experiments use comparable dataset sizes (via subsampling) to ensure fair within-condition comparisons. Slight variations across modules reflect differing experimental goals. This setup supports a systematic evaluation of the data characteristics that underlie effective analytical reasoning, thereby laying the foundation for our investigation.
4.1 Data Comprehension
Tabular Information.
To evaluate the impact of tabular data visibility, we compare two settings: without (w/o Info) and with (w/ Info) table information. In the w/ Info setting, we provide the necessary table information, such as column names, data types, and sample entries; in w/o Info, only the filename (e.g., data.csv) is provided. This ablation tests whether explicit access to table context improves reasoning. As shown in Table 1, adding table information helps the models in simpler tasks like QRData, showing slight improvements in performance. However, the gains are limited, indicating that the models already handle much of the reasoning without explicit table inputs. For the more complex DiscoveryBench, while the 7B model benefits from the inclusion of table information, the 14B model exhibits a slight drop in accuracy. This decline may be related to the increased input length, which could lead the 14B model to generate longer, less focused outputs. We hypothesize that this affects reasoning coherence, though further analysis is needed to confirm the effect.
Model QRData DiscoveryBench
w/o Info w Info w/o Info w Info
Qwen2.5-7B 6.57 7.54 0.42 1.26
Qwen2.5-14B 15.09 15.82 0.42 0.00
Data Complexity.
To evaluate the model’s ability to maintain focus amid distracting information, we introduce additional tables as irrelevant inputs, which act as semantic noise. This setup requires the model to reason over multiple tables, some of which are irrelevant, simulating scenarios with heightened complexity. Models that effectively filter out irrelevant data while focusing on task-relevant information are considered to exhibit stronger data understanding capabilities. For a fair evaluation, we remove task-specific background descriptions from the input, forcing the model to rely solely on the tabular inputs alone. As shown in Table 2, the inclusion of redundant data increases input complexity but does not lead to a significant decline in overall performance. The 7B model exhibits a modest decline, suggesting it is more sensitive to increased input noise. In contrast, the 14B model maintains stable performance, demonstrating stronger filtering capability and resilience in multi-source settings.
Model QRData DiscoveryBench
w/o Extra w/ Extra w/o Extra w/ Extra
Qwen2.5-7B 37.96 34.55 5.44 4.18
Qwen2.5-14B 52.55 52.07 10.88 12.13
Discussion and Implications.
The minimal performance gains from explicit table input suggest that data comprehension is not the primary bottleneck in data analysis. Even under input noise, models maintain stable performance and demonstrate robust data comprehension abilities—likely internalized during pretraining. This indicates that basic data understanding has already been internalized during pretraining, rendering additional context less beneficial.
@@ 4.1 Data Comprehension
Table 3: Performance comparison across models in multi-turn settings (% accuracy).
@@ 4.2 Code Capability
To investigate the role of code generation in data analysis, we evaluate a diverse set of models with varying training objectives and architectures. Rather than treating code correctness as an end in itself, we examine how well models utilize code as part of a broader reasoning process to achieve task success.
Code Performance.
We analyze how different models utilize code during problem solving. Table 3 provides the overall performance across models, while Table 4 summarizes the average code error rates observed across two datasets.
Our analysis reveals several key findings: (1) Code specialization does not guarantee better performance: Qwen2.5-7B-Coder does not demonstrate a clear advantage over general-purpose models. This suggests that code specialization alone may not translate to better performance in analytical tasks, due to limitations in instruction-following or reasoning generalization. (2) Distillation can lead to functional hallucination: R1-Distill-Qwen-7B, despite being distilled from a large reasoning model, performs poorly, often hallucinating file interpretations rather than generating executable code. (3) Long-context capability does not imply efficient task execution: When comparing Qwen2.5-7B-Instruct-1M and Qwen2.5-7B-Instruct under matched output length constraints, both models exhibit comparable coding capabilities; however, the latter demonstrated superior planning efficiency by completing tasks in fewer interaction rounds.
To further substantiate these findings, we manually sample 354 erroneous responses and categorized the errors using GPT-4o-mini. The categorization is based on the gap between the incorrect responses and the corresponding correct trajectories. As shown in Figure 2, only a small fraction of errors stem from syntactic or semantic code defects (e.g., invalid syntax). The majority of errors stem from higher-level reasoning failures, such as incorrect hypothesis formulation or premature termination. This indicates that planning, rather than coding, is more important.
@@
Discussion and Implications.
Our results indicate that while coding proficiency is necessary, it may not be the primary determinant of success in data analysis. Modern instruction-tuned models already possess sufficient code generation capabilities to handle typical analytical operations. What truly distinguishes stronger agents may be their ability to strategically deploy code—to select appropriate operations, sequence them logically, and interpret outputs for iterative reasoning.
@@ 4.3 Strategic Planning
Data analysis is inherently a planning-intensive process that demands careful coordination of data access, transformation, and reasoning steps. Building on our earlier findings that strategic reasoning plays a decisive role in task success, we further investigate how the reasoning capabilities of LLMs shape their performance in complex analytical workflows. Given that instruction-tuned models (e.g., Qwen2.5-7B-Instruct) demonstrate more coherent planning behavior, we use them as the foundation for all subsequent experiments. Specifically, we systematically evaluate model performance along four key aspects: Interaction Turns, Reasoning Length, Task Complexity, and Problem Diversity.
Interaction Turns.
@@
Figure 3 reports the performance of Qwen2.5-7B and Qwen2.5-14B under different dialogue strategies. To ensure a fair comparison, both models were fine-tuned using the same dataset size across all strategies, which is 1020 here. The results reveal consistent trends across the two models: medium-length interactions generally achieve relatively better performance across both datasets, suggesting they provide an optimal balance between reasoning depth and focus. In contrast, short and long interaction strategies yield slightly lower yet relatively stable results. Notably, the mixed strategy consistently underperforms, likely because variable turn lengths disrupt the model’s ability to learn stable interaction patterns. Given the alignment in trends across model scales and the computational efficiency of the 7B variant, we select Qwen2.5-7B for all subsequent experiments. This enables deeper ablation studies while maintaining analytical rigor.
@@
Table 5: Performance across different training data turn lengths (% accuracy).
To better understand the impact of turn length, we examine the performance of various interaction strategies using the full collected dataset. The results are presented in Table 5, from which we derive the following observations:
(i) Turn length preferences are task-dependent.
@@
(iii) Mixed strategies require principled design.
While combining turn lengths may seem beneficial for diversity, the underperformance of mixed strategies suggests that unstructured variation introduces inconsistency, hindering the model’s ability to learn coherent reasoning patterns. Rather than improving flexibility, random mixing may confuse policy learning. Effective use of diverse interaction styles likely requires intentional design, such as curriculum scheduling or adaptive control, rather than uniform combination.
@@
Reasoning Length.
To investigate whether longer reasoning chains from stronger models improve planning and task success, we augment training samples with intermediate <think> segments generated by DeepSeek-R1. These segments aim to capture intermediate reasoning steps that may scaffold better decision-making during multistep analysis.
We evaluate three settings: (1) Original reasoning, which uses the original training data without modification. These samples include reasoning content, but the reasoning is typically shorter and less explicit compared to the augmented settings; (2) Full reasoning, which replaces the original reasoning with full <think> traces for intermediate turns (excluding the first and final turns); and (3) Summarized reasoning, which substitutes the original reasoning with concise summaries (generated by an LLM) of the full traces. The reasoning is inserted only in the middle turns to avoid hallucinations during initial grounding and to maintain brevity in the final answer generation. Experiments are conducted under varying per-turn token budgets (1024, 2048, and 4096) to simulate different interaction constraints. Figure 4 presents the results. We observe that:
(i) Longer reasoning is not always better.
While longer reasoning chains might seem beneficial, the Full setting consistently underperforms the Original across most configurations, particularly in DiscoveryBench. The sharp decline in performance at the 4096-token level suggests that excessive, unfiltered reasoning can overwhelm the model, either by introducing redundancy, amplifying internal noise, or disrupting attention coherence. In contrast, the Summarized setting (though shorter) consistently matches or surpasses baseline accuracy. By distilling critical inferences and removing irrelevant reasoning steps, it reduces noise and enhances focus, leading to more reliable decision-making. These results indicate that reasoning effectiveness depends less on length and more on information relevance and logical coherence. Well-structured, goal-aligned reasoning often outperforms longer but unfocused alternatives.
@@
Increasing the per-turn token budget can improve performance by enabling better integration of reasoning and task-relevant content, as seen in tasks like QRData. However, these gains are often limited, and in some cases, such as with DiscoveryBench, a larger token budget may even reduce performance by amplifying noise or irrelevant information. This suggests that effectiveness in reasoning-intensive tasks depends less on context size and more on information density and coherence. Simply allocating more tokens without improving content quality offers limited value.
Figure 4: Impact of reasoning length on model performance across token budgets.
@@
Table 6: Performance comparison across different task complexity (% accuracy).
Task Complexity.
To assess how data difficulty affects model reasoning, we classify each example based on the performance of models with varying capacities. Specifically, a task is labeled easy if it can be correctly solved by Qwen2.5-7B, medium if only Qwen2.5-14B can solve it, and hard if it requires DeepSeek-R1 to provide the correct answer. Here, each level contains 733 samples to ensure fair comparison.
We train models using data of varying difficulty levels, with the results summarized in Table 6. On QRData, performance generally improves with training data difficulty, with the medium setting yielding the best results. Combining medium and hard data achieves comparable or better performance across both datasets, indicating that exposure to more complex tasks enhances model generalization on structured analytical problems.
To understand this effect, we analyze the models’ dialogue turns and average response length, as illustrated in Figure 5. The results reveal that: As training data difficulty increases, models shift from multi-turn, feedback-dependent interactions to generating comprehensive answers in fewer steps. This suggests that exposure to harder tasks encourages models to internalize reasoning steps—reducing reliance on iterative refinement and increasing reasoning density per turn. The result is more efficient, self-contained decision-making.
Notably, training on a mix of medium and hard data yields the most compact behavior, with the lowest average interaction turns and shortest responses. However, on DiscoveryBench, models trained on easy data perform better. We attribute this to the need for fine-grained, multi-step computation, where a smaller per-step workload improves reliability. These findings highlight a key trade-off: while complex training fosters reasoning efficiency, simpler strategies may remain valuable for tasks requiring extended exploration.
Figure 5: Impact of training data difficulty on interaction patterns. (a) Average number of response rounds of the model. (b) Average output token length of the model.
Problem Diversity.